I am having a hard time understanding elimination using multiplication and using systems equations - example
Y+2x= -1
Y-4=-2x
I appreciate any help - hn

Answer by Alan3354(69443)

I am having a hard time understanding elimination using multiplication and using systems equations - example
Y+2x= -1
Y-4=-2x
-----------
Put the equations in the same form.
y + 2x = -1
y + 2x = 4
---------
These are inconsistent. y+2x can't equal both -1 & 4.
If you graph them, you get parallel lines.
No intersection, no solution.